OpenStudy (anonymous):

How many milliliters of water must be added to 12 ml of a 60% salt solution in order to produce a 42% salt solution? Round your answer to the nearest tenth. 4.7 ml 5.1 ml 7.6 ml 7.2 ml

OpenStudy (mertsj):

Let x = # of ml of water to add 0(x)+.6(12)=.42(x+12)

OpenStudy (jewlzme17):

How many milliliters of water must be added to 12 ml of a 60% salt solution in order to produce a 42% salt solution? Round your answer to the nearest tenth. ------- Equation: salt + salt = salt 0*x + 0.60*12 = 0.42(x+12) Multiply thru by 100 to get: 0 + 60*12 = 42x + 42*12 42x = 18*12 x = 5.1 ml (amt. of water needed)